Find the dimension and the maximum area of a rectangle if its perimeter is 36.
Answer by srinivas.g(540)

perimeter =36
2(length+width)=36
length+width =36/2
length+width =18
possible alternatives:
length =17 width =1 area =17*1 = 17
length =16 width = 2 area =16*2 = 32
length =15 width =3 area = 15*3 =45
length =14 width = 4 area = 14*4 =56
length =13 width =5 area =13*5 =65
length =12 width =6 area =12*6 =72
length =11 width =7 area =11*7 =77
length =10 width =8 area =10*8 =80
length =9 width =9 area =9*9 =81
so maximum area = 81
dimensions = 9 x 9